OpenStudy (anonymous):

x^2+3/2x=3 solve by completing the square.

OpenStudy (anonymous):

x^2 + 3/2x = 3 x^2+3/2x + 9/16 -9/16 = 3 (x+3/4)^2 = 57/16 (x+3/4) = (+/-)((57^0.5)/4) x = ((57^0.5) - 3) / 4 or x = (-(57^0.5)-3) / 4

jhonyy9 (jhonyy9):

\[x ^{2} +3/2x =3\]

jhonyy9 (jhonyy9):

\[2x ^{2}+3x=6\]

jhonyy9 (jhonyy9):

\[2x ^{2}+3x-6=0\]

jhonyy9 (jhonyy9):

x_1,_2=(-3+=-sqrt(9+48))/4 =(-3+/-sqrt(57))/4